### v2

# Object Fidelity Diffusion for Remote Sensing Image Generation

Abstract
High-precision controllable remote sensing image generation is both meaningful and challenging. Existing diffusion models often produce low-fidelity images due to their inability to adequately capture morphological details, which may affect the robustness and reliability of object detection models. To enhance the accuracy and fidelity of generated objects in remote sensing, this paper proposes Object Fidelity Diffusion (OF-Diff), which effectively improves the fidelity of generated objects. Specifically, we are the first to extract the prior shapes of objects based on the layout for diffusion models in remote sensing. Then, we introduce a dual-branch diffusion model with diffusion consistency loss, which can generate high-fidelity remote sensing images without providing real images during the sampling phase. Furthermore, we introduce DDPO to fine-tune the diffusion process, making the generated remote sensing images more diverse and semantically consistent. Comprehensive experiments demonstrate that OF-Diff outperforms state-of-the-art methods in the remote sensing across key quality metrics. Notably, the performance of several polymorphic and small object classes shows significant improvement. For instance, the mAP increases by 8.3%, 7.7%, and 4.0% for airplanes, ships, and vehicles, respectively.
Figure 1: Four critical failure modes in the State-of-the-Art (SOTA) method ($C$-Diff): (a) Control Leakage: content spills outside the specified layout; (b) Structural Distortion: resulting in malformed object morphology; (c) Dense Generation Collapse: a loss of control in crowded scenes; and (d) Feature-level Mismatch: a distributional drift from real data, visualized by t-SNE. Our OF-Diff (2nd row) effectively resolves these issues.
Introduction
Synthesizing high-fidelity, spatially-controllable remote sensing (RS) images is a critical frontier for overcoming the data limitations that hinder downstream perception tasks like object detection (Yang et al. 2021; Zhang et al. 2020; Yang et al. 2019). Current RS generation methods, however, typically rely on either ambiguous text prompts (Khanna et al. 2023; Sebaq and ElHelw 2024) or auxiliary conditions like semantic maps (Sebaq and ElHelw 2024; Tang et al. 2024; Gong et al. 2024; Hu et al. 2025; Jia et al. 2025). While visually plausible, such guidance is fundamentally disconnected from the instance-level ground truth, failing to provide the precise control necessary for effective data augmentation.
In contrast, Layout-to-Image (L2I) generation conditioned on object bounding boxes presents a more robust solution for precise spatial control. While this approach has been extensively developed in the natural image domain with methods like LayoutDiffusion (Zheng et al. 2023) treating it as a multi-modal fusion task, GLIGEN (Li et al. 2023) enabling open-world generation with extra controls, and ODGen (Zhu et al. 2024) decoupling objects for higher controllability, its direct application to RS is non-trivial, due to the unique challenges of RS imagery, such as vast backgrounds, arbitrary object orientations, and dense scenes.
Figure 2: Comparison of OF-Diff with mainstream L2I methods. FG/BG stands for foreground/background. (a) Layout-conditioned baseline. (b) Added instance-based module, limited by quality/quantity of patches from ground truth. (c) OF-Diff enhances fidelity via shape extraction and DDPO, without patch reliance. (d) Results demonstrate superiority.
In RS layout-to-image generation, existing methods like AeroGen (Tang et al. 2025) and CC-Diff (Zhang et al. 2024) take different approaches. AeroGen, a coarse layout-conditioned model, suffers from limited spatial and shape control. In contrast, instance-level methods like CC-Diff achieve higher controllability and fidelity by referencing real instances, but this creates heavy dependence on the quality and quantity of real data, limiting generalization and flexibility. We summarize common failure modes (see Figure 1), including control leakage, structural distortion, dense generation collapse, and feature-level mismatch.
These deficiencies significantly degrade the performance on object detection tasks, limiting their practical application in intelligent RS interpretation. In this paper, we introduce Object Fidelity Diffusion Model (OF-Diff). It is designed to improve the shape fidelity and layout consistency of object generation in RS images. As shown in Figure 2, the existing L2I methods are mainly divided into two categories. The first is layout-conditioned baseline, as shown in Figure 2(a), like Aerogen and LayoutDiffusion. The second is the method with instance-based module, as shown in Figure 2(b), like CC-Diff. However, such methods require real instances and images as references during the sampling stage in order to generate high-quality synthetic images. In contrast, OF-Diff only needs to extract the shape of the foreground to generate high-fidelity RS images, as shown in Figure 2(c). In addition, it fine-tunes the diffusion with DDPO, effectively enhancing the performance of downstream tasks for the generated images. The results in Figure 2(d) demonstrates the superiority of OF-Diff over other methods. Our contributions are summarized as follows:
  • We introduce OF-Diff, a dual-branch controllable diffusion model with prior shape extraction, which improves generation fidelity while reducing reliance on real images, enhancing practical applicability.
  • We propose a controllable generation pipeline that fine-tunes diffusion models with DDPO for remote sensing images, further boosting fidelity and diversity.
  • Extensive experiments demonstrate that OF-Diff generates high-fidelity, layout- and shape-consistent images with dense objects, and serves as an effective enhancement for object detection tasks.
Figure 3: OF-Diff’s overall architecture. (a) During the training phase, the shape features of objects obtained through ESGM and image are feature extracted by Control, and the parameters are updated to the dual-branch stable diffusion. Noisy image $Z_{t}$ in the latent space is processed through shape feature stable diffusion and Mix feature stable diffusion, generating the noise predictions $\varepsilon_{\theta}^{s}$ and $\varepsilon_{\theta}^{m}$. The diffusion consistency loss is calculated by $\varepsilon_{\theta}^{s}$ and $\varepsilon_{\theta}^{m}$, and the denoising loss is calculated by $\varepsilon_{\theta}^{s}$ and $\varepsilon$. (b) During the sampling phase, only label and shape feature stable Diffusion are utilized to generate synthetic images. (c) The pipeline of OF-Diff. DDPO is optional if the generated data is expected to better conform to the distribution of a specific dataset.
Related Work
Advances in Image Generation
[…]
Layout-to-Image Generation
Controllable image synthesis primarily includes text-to-image (T2I) and layout-to-image (L2I) generation. While T2I models (Nichol et al. 2022; Ramesh et al. 2022) achieve semantic alignment via textual prompts, L2I methods offer better spatial control. Recent works enhance layout conditioning through layout-as-modality designs (Zheng et al. 2023), gated attention (Li et al. 2023), and instance-wise generation (Wang et al. 2024; Zhou et al. 2024). However, these methods rely solely on coarse layout inputs (e.g., bounding boxes), which lack fine-grained shape information critical for synthesizing morphologically complex objects.
Remote Sensing Image Synthesis
Synthesizing high-fidelity training data is crucial for advancing remote sensing (RS) object detection, a field critical to numerous applications, but often hampered by the scarcity of extensively annotated datasets. Despite its necessity, most generative models for RS imagery, such as DiffusionSat (Khanna et al. 2023) and RSDiff (Sebaq and ElHelw 2024), still rely on coarse semantic guidance. While other approaches leverage diverse control signals (Tang et al. 2024) like OpenStreetMaps (Espinosa and Crowley 2023), they are generally not optimized for the bounding box format central to object detection. This naturally motivates L2I approaches including AeroGen (Tang et al. 2025) and CC-Diff (Zhang et al. 2024), which have improved spatial accuracy and contextual consistency through layout-mask attention and FG/BG dual re-samplers. However, they suffer from limited controllability and heavy reliance on real data.
Method
Preliminary
Diffusion models (Song, Meng, and Ermon 2021) aim to capture the underlying data distribution $p(x)$ by iteratively reconstructing data from a noisy representation that is initially sampled from a standard normal distribution.
Denoising Diffusion Probabilistic Models (Ho, Jain, and Abbeel 2020) parameterize the model as the function $\epsilon_{\theta}(x_{t},t)$ to predict the noise component of the sample $x_{t}$ at any time step $t$. The training objective is to minimize the mean squared error (MSE) loss between the actual noise $\epsilon$ and the predicted noise $\epsilon_{\theta}(x_{t},t)$:
$\mathcal{L}=\mathbb{E}_{x_{t},t,\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\left[\left\|\epsilon_{\theta}\left(x_{t},t\right)-\epsilon\right\|^{2}\right].$ (1)
Stable Diffusion (Rombach et al. 2022a; Qiu et al. 2025) utilizes a pre-trained VQ-VAE (Van Den Oord, Vinyals et al. 2017) to encode images into a lower-dimensional latent space, performing training on the latent representation $z_{0}$. In the context of conditional generation, given a text prompt $c_{t}$ and task-specific conditions $c_{f}$, the diffusion training loss at time step $t$ can be expressed as:
$\mathcal{L}=\mathbb{E}_{z_{t},t,c_{t},c_{f},\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\left[\left\|\epsilon-\epsilon_{\theta}\left(z_{t},t,c_{t},c_{f}\right)\right\|^{2}\right].$ (2)
where $\mathcal{L}$ represents the overall learning objective of the complete diffusion model. This objective function is explicitly applied during the fine-tuning of diffusion models in conjunction with ControlNet (Zhang, Rao, and Agrawala 2023).
Architecture of OF-Diff
As illustrated in Figure 3(a), during the training phase of OF-Diff, both the real images and their corresponding labels are needed for diffusion training. First, they are fed into the Enhanced Shape Generation Module (ESGM) for shape extraction, yielding the object mask. Subsequently, the real image and its object mask are fed into ControlNet for feature extraction, resulting in the image feature $c_{i}$ and the label feature $c_{l}$. Second, the input image is first compressed into the latent space features $z_{0}$ by a pre-trained VQ-VAE. Then, it is concatenated with random Gaussian noise $\epsilon$ to form $z_{t}$. After being injected into the stable diffusion encoder blocks, the feature $Z_{t}$ is fed into a dual-branch decoder architecture. The two components of this architecture are respectively designated as the shape feature stable diffusion decoder and the mix feature stable diffusion decoder. Third, a weighted summation of $c_{i}$ and the stop - gradient $c_{l}$ is carried out to $c_{m}$. $c_{m}$ is defined as:
$c_{\operatorname{m}}=\frac{n}{N}\cdot c_{i}+\operatorname{sg}\left[c_{l}\right],$ (3)
where $n$ denotes the current iteration number, and $N$ is the total number of iterations. In order to enable the prediction from the mix feature stable diffusion to serve as an anchor point, guiding the convergence trajectory of the shape stable diffusion and thus improving the morphological fidelity of the image, we adopt a stop-gradient strategy (Chen and He 2021) for $c_{l}$ when calculating $c_{m}$. The loss $\mathcal{L}_{s}$ and $\mathcal{L}_{m}$ of shape feature stable diffusion and mix feature stable diffusion are defined as:
$\mathcal{L}_{s}=\mathbb{E}\left[\left\|\epsilon_{\theta}^{s}-\epsilon\right\|^{2}\right],\epsilon_{\theta}^{s}=\epsilon_{\theta}\left(z_{t},t,c_{t},c_{s}\right),$ (4)
$\mathcal{L}_{m}=\mathbb{E}\left[\left\|\epsilon_{\theta}^{m}-\epsilon\right\|^{2}\right],\epsilon_{\theta}^{m}=\epsilon_{\theta}\left(z_{t},t,c_{t},c_{m}\right),$ (5)
To assist the shape feature stable diffusion in breaking away from local minima with low fidelity and migrating towards regions of high fidelity in the parameter space, we introduce a diffusion consistency loss:
$\mathcal{L}_{c}=\mathbb{E}\left[\left\|\epsilon_{\theta}^{m}-\operatorname{sg}\left[\epsilon_{\theta^{\prime}}^{\operatorname{mix}}\right]\right\|^{2}\right],$ (6)
The predicted noise $\epsilon_{\theta}^{m}$ from the mix feature stable diffusion is more accurate than $\epsilon_{\theta}^{s}$ predicted by the shape feature stable diffusion, owing to the additional image prior control. Nevertheless, during the sampling process, we do not utilize inputs with real images. Consequently, we employ a stop-gradient operation so that the more accurate $\epsilon_{\theta}^{m}$ can serve as an anchor, guiding the convergence trajectory of the masked diffusion towards high-fidelity local minima in the parameter space. Ultimately, the loss employed to train OF-Diff is defined as:
$\mathcal{L}=\mathcal{L}_{s}+\mathcal{L}_{m}+\mathcal{L}_{c},$ (7)
During the sampling phase, as illustrated in Figure 3(b), only the frozen ControlNet and the shape feature stable diffusion are utilized with arbitrary label prior control to synthesize RS images.
Figure 4: The architecture of Enhanced Shape Generation Module (ESGM).
Enhanced Shape Generation Module
In natural imagery, perspective and scale changes prevent a unique geometric model for most objects. Conversely, remote-sensing objects display quasi-invariant shapes. For instance, courts are rectangular, chimneys and oil tanks circular, and airplanes bilaterally symmetric with a distinct nose and tail. This shape consistency enables the use of masks to impose strong controllability on image synthesis for remote sensing. To better leverage category labels for object shape extraction, we introduce the Enhanced Shape Generation Module (ESGM, see Figure 4). During the training phase, ESGM uses paired images and labels to generate precise object masks. And at sampling time, it employs learned shape priors to synthesize diverse masks of object shape.
For the given image $x_{i}$ and its bounding box $y_{i}^{j}$ corresponding to category $j$ $(j\in[1,N])$, we first utilize the RemoteCLIP (Liu et al. 2024) to generate a textual description of the object enclosed within the bounding box. With this description and the original image $x_{i}$, the RemoteSAM (Yao et al. 2025) then generates the corresponding shape mask $x_{i}^{j}$.
In the shape augmentation phase, we crop $x_{i}^{j}$ using its bounding box $y_{i}^{j}$ to extract a shape mask $x_{i}^{j}$ that contains only the target object. Subsequently, we apply a random rotation to $x_{i}^{j}$ and reposition the rotated mask onto a blank canvas using the original bounding box coordinates $y_{i}^{j}$, resulting in a shape-enhanced mask. During the training phase, ESGM can directly obtain the shape of real images. While sampling, ESGM selects enhanced shapes from a mask pool collected during or beyond training, as in common data-efficient practices where segmentation masks are reused or synthesized with minimal impact on generalization. In our setup, we use masks generated during training.
  Method DIOR Dataset DOTA Dataset
11
[…]
LayoutDiff 37.60 0.015 0.447 70.32 7.01 21.73 0.015 0.288 77.56 21.43
GLIGEN 35.06 0.010 0.622 76.41 6.51 39.79 0.026 0.357 76.19 15.58
AeroGen 27.78 0.013 0.563 81.69 55.38 26.65 0.017 0.298 81.91 44.85
$C$-Diff 49.62 0.024 0.685 82.61 42.17 32.40 0.019 0.279 81.63 49.62
Ours 24.92 0.011 0.312 82.55 58.99 20.84 0.014 0.271 83.79 55.68
Table 1: Quantitative comparison with SOTA methods on DIOR and DOTA. We evaluate performance on generation fidelity (FID, KID, CMMD) and layout consistency (CAS, YOLOScore). OF-Diff demonstrates superior overall performance. The best and second best results are highlighted. Lower is better ($\downarrow$), higher is better ($\uparrow$).
  Method DIOR Dataset DOTA Dataset
7
mAP mAP50 mAP75 mAP mAP50 mAP75
Baseline 30.51 52.84 32.10 38.09 66.31 38.44
LayoutDiff 29.81 52.14 30.36 38.91 66.75 40.37
GLIGEN 28.48 51.27 29.21 38.84 66.10 40.24
AeroGen 31.53 53.37 33.60 38.45 67.09 39.07
$C$-Diff 31.82 53.48 33.97 38.51 66.52 39.02
Ours 32.71 54.44 34.05 40.03 67.89 42.20
Table 2: Trainability ($\uparrow$) comparison on DIOR and DOTA. ‘Baseline’ denotes accuracy with the unaugmented dataset.
DDPO fine-tuning
To enhance the diversity of the distribution of data generated by the fine-tuned model and maintain better consistency with the distribution of real images (Schulman et al. 2015, 2017), denoising diffusion policy optimization (DDPO) (Black et al. 2023) is applied in the post-training of OF-Diff. DDPO regards the denoising process of the diffusion model as a multi-step Markov decision process (MDP). The mapping relationship is defined as follows:
$\pi(a_{t}\mid s_{t})\triangleq p_{\theta}(\mathbf{x}_{t-1}\mid\mathbf{x}_{t},c)$ (8)
$P(s_{t+1}\mid s_{t},a_{t})\triangleq(\delta_{c},\,\delta_{t-1},\,\delta_{\mathbf{x}_{t-1}})$ (9)
$\rho_{0}(s_{0})\triangleq\bigl{(}p(c),\,\delta_{T},\,\mathcal{N}(0,I)\bigr{)}$ (10)
$R(s_{t},a_{t})\triangleq\begin{cases}r(\mathbf{x}_{0},c),&\text{if }t=0,\\ 0,&\text{otherwise}.\end{cases}$ (11)
where $\delta_{z}$ denotes the Dirac delta distribution whose probability density is zero everywhere except at $z$. The symbols $s_{t}$ and $a_{t}$ represent the state and action at time $t$, respectively. Specifically, $s_{t}$ is defined as the tuple composed of the condition $c$, the time step $t$, and the noisy image $\mathbf{x}_{t}$ at that time, whereas $a_{t}$ is defined as the noisy image $\mathbf{x}_{t-1}$ from the preceding time step. The policy is denoted by $\pi(a_{t}\mid s_{t})$, the transition kernel by $P(s_{t+1}\mid s_{t},a_{t})$, the initial state distribution by $\rho_{0}(s_{0})$, and the reward function by $R(s_{t},a_{t})$. To optimize the policy $\pi(a_{t}\mid s_{t})$ so as to maximize the cumulative reward $\mathbb{E}_{\tau\sim p(,\cdot\mid\pi)}!\Bigl{[}\sum_{t=0}^{T}R(\mathbf{s}_{t},\mathbf{a}_{t})\Bigr{]}$, the gradient $\hat{g}$ is computed as follows:
$\hat{g}$= $\mathbb{E}\!\biggl{[}\sum_{t=0}^{T}\frac{p_{\theta}(\mathbf{x}_{t-1}\mid c,\ t,\ \mathbf{x}_{t})}{p_{\theta^{\prime}}(\mathbf{x}_{t-1}\mid c,\ t,\ \mathbf{x}_{t})}$ (12)
$\cdot r(\mathbf{x}_{0},c)\cdot\nabla_{\theta}\log p_{\theta}(\mathbf{x}_{t-1}\mid c,\ t,\ \mathbf{x}_{t})\biggr{]}$
$r(\mathbf{x}_{0},c)=\bigl{(}KNN(\mathbf{x}_{0},\mathbf{x}_{0})-\omega KL(\mathbf{x}_{0},\mathbf{x}_{0}^{\prime})\bigr{)}$ (13)
Figure 5: Qualitative results on DIOR and DOTA. OF-Diff is more realistic and fidelity compared to other methods.
The reward functions based on K-Nearest Neighbor (KNN) and KL divergence are introduced, respectively, to optimize the diversity of generated data and the distribution consistency between generated data and real data. $\omega$ is the weight parameter, and $\mathbf{x}_{0}^{\prime}$ is the real image in the dataset.
Experiments
Experimental Settings
Datasets.
  • DIOR-R (Cheng et al. 2022) is a rotated version of the DIOR benchmark (Li et al. 2020), annotated with oriented bounding boxes, 20 categories. We follow the official split, which divides the dataset into training, validation, and test sets with an approximate ratio of 1:1:2. All experiments are trained on the training set and evaluated on the test set.
  • DOTA-v1.0 (Xia et al. 2018) includes 15 common object categories, with 2,806 images and 188,282 instances. This dataset is more challenging due to the presence of numerous dense scenes and small objects. Following MMRotate (Zhou et al. 2022), each image in the training and test sets is cropped into $512\times 512$ tiles, and those without any valid objects are removed. All experiments are conducted on these processed images.
Implementation Details. We train OF-Diff separately on each dataset (DIOR/DOTA), based on the Stable Diffusion 1.5 (Rombach et al. 2022b) pretrained model. Only the ControlNet and shape feature SD decoder are fine-tuned, while all other modules remain frozen. Training is performed using the AdamW optimizer with a learning rate of 1e-5. The global batch size is set to 64, and training runs for 100 epochs.
Benchmark Methods. We compare our method with state-of-the-art L2I generation models for both remote sensing (AeroGen (Tang et al. 2025), CC-Diff (Zhang et al. 2024)), and natural images (LayoutDiffusion (Zheng et al. 2023), GLIGEN (Li et al. 2023)). For a fair comparison, all models are re-trained using our dataset settings, following their official training details respectively.
Evaluation Metrics. To more comprehensively evaluate the effectiveness of OF-Diff, we adopt a total of 13 metrics spanning 4 different evaluation aspects.
  • Generation Fidelity. We use FID (Heusel et al. 2017) and KID (Bińkowski et al. 2018) to assess perceptual quality, along with CMMD (Jayasumana et al. 2024), which measures CLIP feature distances between generated and real images to evaluate layout alignment.
  • Layout Consistency. We report CAS (Ravuri and Vinyals 2019) using a pretrained classifier to assess object recognizability, and YOLOScore by applying a pretrained Oriented R-CNN (Xie et al. 2021) (w/. Swin Transformer backbone (Liu et al. 2021), MMRotate) to generated images for instance-level consistency.
  • Shape Fidelity. To assess the geometric quality of generated instances, we perform pairwise comparisons with ground-truth shapes. Each instance pair is cropped, resized to $\times64$, and converted to edge maps. We compute five metrics: IoU, Dice, Chamfer Distance (CD), Hausdorff Distance (HD), and SSIM (Wang et al. 2004).
[…]
Qualitative Results
Comparative Results. Figure 5 compares the generation results of OF-Diff with other methods. OF-Diff not only generates more realistic images but also has the best controllability. For instance, in the first two cases, OF-Diff successfully controlled the number and layout information of the generated objects. The third and fourth cases demonstrate the accuracy of OF-Diff in generating small targets, which other algorithms fail to do accurately. The last case shows the superiority of OF-Diff over other algorithms when generating objects with complex shapes such as airplane.
Diversity Results. Figure 6 shows that the images generated by OF-Diff consistently present plausible textures and realistic object shapes. For instance, airplanes rendered at different orientations maintain coherent semantic relationships with their surrounding environments. Even in small-object scenes (some of which are grayscale remote-sensing images from the DOTA dataset), OF-Diff can still generate visually faithful and geometrically accurate results.
  Method Dataset IoU $\uparrow$ Dice $\uparrow$ CD $\downarrow$ HD $\downarrow$ SSIM $\uparrow$
LayoutDiff 0.0497 0.0908 12.037 25.962 0.1667
GLIGEN 0.055 0.1002 12.257 25.850 0.1652
AeroGen DIOR 0.0855 0.153 8.209 20.314 0.2142
$C$-Diff 0.0891 0.1582 8.0909 20.066 0.1963
Ours 0.1009 0.1763 7.6579 19.459 0.2691
LayoutDiff 0.0402 0.0748 15.229 30.202 0.2194
GLIGEN 0.0645 0.1182 10.432 23.196 0.1967
AeroGen DOTA 0.0863 0.1536 8.1386 20.687 0.2261
$C$-Diff 0.0692 0.1255 9.6226 21.247 0.2171
Ours 0.1205 0.2045 6.6317 17.311 0.2938
Table 3: Object-Shape Fidelity on Canny Edge Maps. We measure the morphological similarity between generated and ground-truth instances by computing IoU, DICE, Chamfer Distance (CD), Hausdorff Distance (HD), and SSIM.
  Method Unknown Layout during Training
7
[…]
YOLO
[…]
LayoutDiff 44.58 0.018 0.539 29.34 10.37 53.07
GLIGEN 39.56 0.013 0.444 66.36 2.13 52.68
AeroGen 28.62 0.013 0.276 80.78 46.36 55.11
$C$-Diff 49.92 0.024 0.513 78.01 51.74 53.72
Ours 24.18 0.012 0.271 83.34 49.59 56.65
Table 4: Quantitative comparison on the unknown layout dataset during training (DIOR Val).
Figure 6: The diversity of different results from the same OF-Diff model.
Figure 7: AP50 on DIOR and DOTA.
Quantitative Results
Generation Fidelity and Consistency. We compared OF-Diff with state-of-the-art generation methods in remote sensing, including layoutDiffusion (Zheng et al. 2023), GLIGEN (Li et al. 2023), AeroGen (Tang et al. 2025), and CC-Diff (Zhang et al. 2024). The performance of these methods is reported in Table 1. OF-Diff achieved nearly the best performance in both generation fidelity metrics (FID, KID, CMMD) and layout consistency metrics, especially on the DOTA dataset.
Trainability of Object Detection. Following the data enhancement protocol in (Chen et al. 2023), we double the training samples using OF-Diff and assess detection results with the expanded dataset. As shown in Table 2, OF-Diff performs the best on both DIOR and DOTA. Specifically, OF-Diff improves mAP by 2.2% on DIOR and 1.94% on DOTA compared to baseline. Notably, the performance of several polymorphic and small object classes shows significant improvement. According to Figure 7, the AP50 increases by 8.3%, 7.7%, and 4.0% for airplane, ship, and vehicle on DIOR, and 7.1%, 5.9% and 4.4% for swimming pool, small vehicle, and large vehicle on DOTA.
Object-Shape Fidelity. We measure the morphological similarity between the generated instances and ground truth by calculating the Intersection over Union (IoU), DICE coefficient, Chamfer distance (CD), Hausdorff distance (HD), and Structural Similarity Index (SSIM), based on the Canny Edge Map. As shown in Table 3, the results demonstrate that OF-Diff attains state-of-the-art performance in all evaluation metrics for object-shape fidelity. Specifically, we first convert the rotated bounding box (R-Box) to a horizontal bounding box (H-Box) and crop the instance with a 20% padding to ensure the full object is captured. The cropped patches are then resized to $64\times 64$ pixels, and their shapes are extracted using cv2.Canny. For a detailed qualitative comparison, Figure 9 visualizes the instance patches and their corresponding edge maps from different methods. Each image set is ordered as follows: Ground Truth, OF-Diff, AeroGen, CC-Diff, GLIGEN, and LayoutDiff, demonstrating our method’s superior ability to adhere to object shapes.
Adaptability of Unknown Layout. To better demonstrate the scalability of these methods, we also generate images based on the unknown layouts during the training phase. According to Table 4, for unknown layout, OF-Diff performs well in terms of generation fidelity, layout consistency, and trainability. In downstream tasks, OF-Diff still delivers a 1.54% mAP gain over the second-best method.
The Detailed Results of Downstream. Table 6 and  7 report the average precision (AP) obtained by the competing generative methods over multiple categories in the downstream tasks. From Tables 6, it can be observed that OF-Diff (ours) achieves a clear advantage in several categories. For instance, OF-Diff achieves superior performance on Airplane (71.3%), Golf Field (75.4%) and Ship (70.5%), with improvements of approximately 5% to 10% over the second-best method. For a few other categories, OF-Diff does not deliver the top AP, yet the gap to the best result remains marginal. Table 7 shows that, on the DOTA dataset, OF-Diff obtains the highest AP in roughly half of all categories and still delivers notable gains in categories such as Small-vehicle (68.3%), Ship (84.4%) and Swimming-pool (67.9%).
ESGM DCLoss DDPO CAS $\uparrow$
YOLOScore $\uparrow$
mAP50 $\uparrow$
80.27 41.20 52.13
82.16 55.08 52.76
81.57 46.27 53.14
81.91 47.74 53.21
82.30 57.83 54.31
82.55 58.99 54.44
Table 5: Ablation study: impact of ESGM, DCLoss, and DDPO on semantic consistency (CAS) and downstream trainability (YOLOScore and mAP50).
Figure 8: The influence of caption on the generation of images in terms of being more realistic and more aesthetically pleasing.
Ablation Study
We assessed the impact of different modules on image generation semantic consistency and downstream trainability by OF-Diff in Table 5. We found that the input of captions would affect the fidelity of image generation. Images generated with captions are more in line with semantic consistency and human aesthetics, but the fidelity of these images decreases. This is equivalent to the data distribution deviating from the real dataset and being more inclined towards the data distribution during pre-training. Therefore, the ablation experiments for each module were conducted based on the absence of caption input. The contribution of each module to the enhancement of image generation fidelity is evaluated by incorporating additional components into the diffusion model with dual-branch. DDPO indicates whether to fine-tune the trained diffusion model through reinforcement learning. Results show that Enhanced Shape Generation Module (ESGM), Diffusion Consistency Loss (DCLoss) and the DDPO based on KNN and KL Divergence effectively improve the performance metrics. Notably, ESGM can substantially improve the YOLOScore by over 10%.
Discussion
As shown in Figure 8, the inclusion of additional captions as input has a significant impact on the outcomes of image generation. Specifically, incorporating captions enhances the aesthetic appeal of the generated images, resulting in richer and more visually pleasing color compositions. However, this improvement comes at a cost: similar to CC-Diff, it leads to a deviation of the generated data distribution from that of the original real data. In contrast, when no additional captions are provided as input, although the generated images may appear less aesthetically refined, their data distribution remains closer to that of real images. Furthermore, experimental results demonstrate that this approach yields superior performance in terms of image quality metrics and downstream task effectiveness.
According to the current experimental results, adding the DDPO strategy does not simultaneously outperform previous results on all metrics. Using reinforcement learning strategies can indeed improve the performance of downstream tasks, but it does not necessarily improve the quality of image generation simultaneously. In other words, reinforcement learning strategies can also purposefully improve the quality of image generation, but this may come at the cost of not improving the performance of downstream tasks.
Limitations
The proposed OF-Diff injects object shape masks extracted from the image layout as controllable conditions into the diffusion model, which effectively enhances object fidelity and improves the generation of small objects. However, this also makes the model heavily dependent on the quality of the extracted shape masks. If the obtained masks are already distorted, the resulting generated images will likewise fail to achieve satisfactory quality.
Conclusion
Existing image generation methods struggle to precisely generate dense small objects and those with complex shapes, such as numerous small vehicles and airplanes in remote sensing images. To address this, we introduce OF-Diff, a dual-branch controllable diffusion model with prior shapes extraction and DDPO. During the training phase, we extract the prior shape of objects to enhance controllability and use a dual-branch diffusion with parameter sharing to improve the model’s learning ability for real images. Therefore, in the sampling phase, OF-Diff can generate images with high fidelity without real images as references. Finally, we fine-tune the diffusion by DDPO that combines KNN and KL divergence to make the synthesized images more realistic and consistent. Extensive experiments demonstrate the effectiveness and superiority of OF-Diff in generating small and difficult objects with complex structures and dense scenes in remote sensing.