### v3

# Object Fidelity Diffusion for Remote Sensing Image Generation

Abstract
High-fidelity, controllable remote sensing layout-to-image generation is highly valuable for providing high-quality data for downstream object detection tasks. However, existing methods either rely on additional textual guidance, leading to geometric distortions, or require extra real-image references, limiting practical applicability. To address these challenges, we propose Object Fidelity Diffusion (OF-Diff), which leverages object layouts to extract structural shape priors and employs an online-distillation strategy to integrate complex image features. This allows the model to perform highly controllable, high-fidelity image generation at inference without relying on real-image references. Furthermore, we introduce DDPO to fine-tune the diffusion process, making the generated remote sensing images more diverse and semantically consistent. Comprehensive experiments demonstrate that OF-Diff outperforms state-of-the-art methods in the remote sensing across key quality metrics. Notably, the performance of several polymorphic and small object classes shows significant improvement. For instance, the mAP increases by 8.3%, 7.7%, and 4.0% for airplanes, ships, and vehicles, respectively.
[…]
1 Introduction
Figure 1: Four critical failure modes in the State-of-the-Art (SOTA) method ($C$-Diff): a distributional drift from real data, visualized by t-SNE; and (a) Control Leakage; (b) Structural Distortion; (c) Dense Generation Collapse. Our OF-Diff (2nd row) effectively resolves these issues.
Synthesizing high-fidelity, spatially-controllable remote sensing (RS) images is a critical frontier for overcoming the data limitations that hinder downstream perception tasks like object detection [yang2021r3det, zhang2020well, Yang_2019_ICCV]. Current RS generation methods, however, typically rely on either ambiguous text prompts [khanna2023diffusionsat, sebaq2024rsdiff] or auxiliary conditions like semantic maps [sebaq2024rsdiff, tang2024crs, gong2024crossearth, hu2025earth, jia2025can]. While visually plausible, such guidance is fundamentally disconnected from the instance-level ground truth, failing to provide the precise control necessary for effective data augmentation.
In contrast, Layout-to-Image (L2I) generation conditioned on object bounding boxes offers a more robust solution for precise spatial control. This paradigm has been extensively studied in the natural image domain—LayoutDiffusion [zheng2023layoutdiffusion] treats it as a multi-modal fusion problem, GLIGEN [li2023gligen] enables open-world generation through additional control signals, and ODGen [zhu2024odgen] improves controllability by decoupling objects—yet its direct application to remote sensing (RS) imagery remains non-trivial due to expansive backgrounds, arbitrary object orientations, and densely packed scenes.
In RS layout-to-image generation, existing methods like AeroGen [tang2025aerogen] and CC-Diff [zhang2024cc] take different approaches. AeroGen, a coarse layout-conditioned model, suffers from limited spatial and shape control. In contrast, instance-level methods like CC-Diff achieve higher controllability and fidelity by referencing real instances, but this creates heavy dependence on the quality and quantity of real data, limiting generalization and flexibility.The images generated via CC-Diff diverge more markedly from the real remote sensing data distribution, aligning instead with the style characteristic of the model’s pre-training corpus. We summarize common failure modes (see Figure 1), including control leakage, structural distortion, dense generation collapse and feature-level mismatch.
These deficiencies significantly degrade the performance on object detection tasks, limiting their practical application in intelligent RS interpretation. In this paper, we introduce Object Fidelity Diffusion Model (OF-Diff). It is designed to improve the shape fidelity and layout consistency of object generation in RS images. As shown in Figure 2, the existing L2I methods are mainly divided into two categories. The first is layout-conditioned baseline, as shown in Figure 2(a), like Aerogen and LayoutDiffusion. The second is the method with instance-based module, as shown in Figure 2(b), like CC-Diff. However, such methods require real instances and images as references during the sampling stage in order to generate high-quality synthetic images. In contrast, OF-Diff generates high-fidelity remote-sensing objects using only the foreground shape, and subsequently employs online-distillation to further align the outputs with real images, as shown in Figure 2(c). In addition, it fine-tunes the diffusion with DDPO, effectively enhancing the performance of downstream tasks for the generated images. The results in Figure 2(d) demonstrates the superiority of OF-Diff over other methods. Our contributions are summarized as follows:
  • We introduce OF-Diff, an online-distillation controllable diffusion model with prior shape extraction, which improves generation fidelity while reducing reliance on real images, enhancing practical applicability.
  • We propose a controllable generation pipeline that fine-tunes diffusion models with DDPO for remote sensing images, further boosting fidelity and diversity.
  • Extensive experiments demonstrate that OF-Diff generates high-fidelity, layout- and shape-consistent images with dense objects, and serves as an effective enhancement for object detection tasks.
Figure 2: Comparison of OF-Diff with mainstream L2I methods. FG/BG stands for foreground/background. (a) Layout-conditioned baseline. (b) Added instance-based module, limited by quality/quantity of patches from ground truth. (c) OF-Diff enhances fidelity via shape extraction and DDPO, without patch reliance. (d) Results demonstrate superiority.
2 Related Work
2.1 Advances in Image Generation
[…]
2.2 Layout-to-Image Generation
Controllable image synthesis primarily includes text-to-image (T2I) and layout-to-image (L2I) generation. While T2I models [nichol2022glide, ramesh2022hierarchical] achieve semantic alignment via textual prompts, L2I methods offer better spatial control. Recent works enhance layout conditioning through layout-as-modality designs [zheng2023layoutdiffusion], gated attention [li2023gligen], and instance-wise generation [wang2024instancediffusion, Zhou_2024_CVPR]. However, these methods rely solely on coarse layout inputs (e.g., bounding boxes), which lack fine-grained shape information critical for synthesizing morphologically complex objects.
2.3 Remote Sensing Image Synthesis
Synthesizing high-fidelity training data is crucial for advancing remote sensing (RS) object detection, a field critical to numerous applications, but often hampered by the scarcity of extensively annotated datasets. Despite its necessity, most generative models for RS imagery, such as DiffusionSat [khanna2023diffusionsat] and RSDiff [sebaq2024rsdiff], still rely on coarse semantic guidance. While other approaches leverage diverse control signals [tang2024crs] like OpenStreetMaps [espinosa2023generate], they are generally not optimized for the bounding box format central to object detection. This naturally motivates L2I approaches including AeroGen [tang2025aerogen] and CC-Diff [zhang2024cc], which have improved spatial accuracy and contextual consistency through layout-mask attention and FG/BG dual re-samplers. However, they suffer from limited controllability and heavy reliance on real data.
3 Method
3.1 Preliminary
Diffusion models [song2021denoising] aim to capture the underlying data distribution $p(x)$ by iteratively reconstructing data from a noisy representation that is initially sampled from a standard normal distribution. Denoising Diffusion Probabilistic Models [ho2020denoising] parameterize the model as the function $\epsilon_{\theta}(x_{t},t)$ to predict the noise component of the sample $x_{t}$ at any time step $t$. The training objective is to minimize the mean squared error (MSE) loss between the actual noise $\epsilon$ and the predicted noise $\epsilon_{\theta}(x_{t},t)$:
$\mathcal{L}=\mathbb{E}_{x_{t},t,\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\left[\left\|\epsilon_{\theta}\left(x_{t},t\right)-\epsilon\right\|^{2}\right].$ (1)
Stable Diffusion (SD) [Rombach_Blattmann_Lorenz_Esser_Ommer_2022, qiu2025noise] utilizes a pre-trained VQ-VAE [van2017neural] to encode images into a lower-dimensional latent space, performing training on the latent representation $z_{0}$. In the context of conditional generation, given a text prompt $c_{t}$ and task-specific conditions $c_{f}$, the diffusion training loss at time step $t$ can be expressed as:
$\mathcal{L}=\mathbb{E}_{z_{t},t,c_{t},c_{f},\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\left[\left\|\epsilon-\epsilon_{\theta}\left(z_{t},t,c_{t},c_{f}\right)\right\|^{2}\right].$ (2)
where $\mathcal{L}$ represents the overall learning objective of the complete diffusion model. This objective function is explicitly applied during the fine-tuning of diffusion models in conjunction with ControlNet [zhang2023adding].
Figure 3: OF-Diff’s overall architecture. (a) During training, object shape features extracted by ESGM and image features are processed by ControlNet, and the resulting information is used to update stable diffusion decoders via online-distillation. (b) During sampling, only the label and the shape feature stable diffusion decoder are used to generate synthetic images. (c) Architecture of the Enhanced Shape Generation Module (ESGM).
3.2 Architecture of OF-Diff
As illustrated in Figure 3(a), the training of OF-Diff requires both real images and their corresponding labels. First, for ControlNet, the real image and its label are processed by the Enhanced Shape Generation Module (ESGM) to extract the object mask. The image and mask are then fed into ControlNet to obtain the image feature $c_{i}$ and the shape feature $c_{s}$. To enrich the structural-only shape prior with richer appearance and contextual cues from the image, we combine them into a mix-feature $c_{m}$, which will later serve as a teacher input in online-distillation. Concretely:
$c_{\operatorname{m}}=\frac{n}{N}\cdot c_{i}+\operatorname{sg}\left[c_{s}\right],$ (3)
where $n$ denotes the current iteration number, and $N$ is the total number of iterations. In order to enable the prediction conditioned on mix-feature to serve as a stable anchor point, to improve the morphological fidelity of the generation, we adopt a stop-gradient strategy [Chen_He_2021] for $c_{s}$ when calculating $c_{m}$.
Second, for Stable Diffusion, the input image is first compressed into latent space features $z_{0}$ by a pre-trained VQ-VAE. Then, it is concatenated with Gaussian noise $\epsilon$ to form $z_{t}$. After passing through the SD encoder blocks, the feature $Z_{t}$ is fed into a dual-decoder architecture. One branch, the shape-feature SD decoder, conditions on $c_{s}$; the other, the mix-feature SD decoder, conditions on $c_{m}$. Their reconstruction losses are defined as $L_{s}$ and $L_{M}$, respectively:
$\mathcal{L}_{s}=\mathbb{E}\left[\left\|\epsilon_{\theta}^{s}-\epsilon\right\|^{2}\right],\epsilon_{\theta}^{s}=\epsilon_{\theta}\left(z_{t},t,c_{t},c_{s}\right),$ (4)
$\mathcal{L}_{m}=\mathbb{E}\left[\left\|\epsilon_{\theta}^{m}-\epsilon\right\|^{2}\right],\epsilon_{\theta}^{m}=\epsilon_{\theta}\left(z_{t},t,c_{t},c_{m}\right),$ (5)
Third, for online distillation, the mix-feature SD produces more accurate predictions thanks to its stronger image prior, but needs real images, limiting diversity. In contrast, the shape-feature SD supports arbitrary label control but risks converging to low-fidelity local minima. To reconcile these trade-offs, we propose an online-distillation framework with a consistency loss $L_{c}$:
$\mathcal{L}_{c}=\mathbb{E}\left[\left\|\epsilon_{\theta}^{s}-\operatorname{sg}\left[\epsilon_{\theta^{\prime}}^{\operatorname{m}}\right]\right\|^{2}\right].$ (6)
Here, the prediction $\epsilon_{\theta^{\prime}}^{m}$ from mix-feature SD decoder acts as a stop-gradient teacher signal, serving as an anchor to guide the prediction $\epsilon_{\theta}^{s}$ from shape-feature SD decoder towards high-fidelity optima in parameter space.
The overall training objective is therefore:
$\mathcal{L}=\mathcal{L}_{s}+\mathcal{L}_{m}+\lambda\mathcal{L}_{c},$ (7)
During the sampling phase, as illustrated in Figure 3(b), only the frozen ControlNet and the shape feature stable diffusion are utilized with arbitrary label prior control to synthesize RS images.
3.3 Enhanced Shape Generation Module
In natural imagery, perspective and scale changes prevent a unique geometric model for most objects. Conversely, remote-sensing objects display quasi-invariant shapes. For instance, courts are rectangular, chimneys and oil tanks circular, and airplanes bilaterally symmetric with a distinct nose and tail. This shape consistency enables the use of masks to impose strong controllability on image synthesis for remote sensing. To better leverage category labels for object shape extraction, we introduce the Enhanced Shape Generation Module (ESGM, see Figure 3(c)). During the training phase, ESGM uses paired images and labels to generate precise object masks. And at sampling time, it employs learned shape priors to synthesize diverse masks of object shape.
For the given image $x_{i}$ and its bounding box $y_{i}^{j}$ corresponding to category $j$ $(j\in[1,N])$, we first utilize the RemoteCLIP [liu2024remoteclip] to generate a textual description of the object enclosed within the bounding box. With this description and the original image $x_{i}$, the RemoteSAM [yao2025remotesam] then generates the corresponding shape masks $x_{i}^{j}$}.
In the shape augmentation phase, each object mask $x_{i}^{j}$ is cropped by its bounding box $y_{i}^{j}$, randomly rotated, and placed back onto a blank canvas to produce a shape-enhanced mask. During training, ESGM uses real image shapes; at sampling, it selects enhanced shapes from a lightweight mask pool collected during or after training. In our experiments, we use masks generated during training.
3.4 DDPO fine-tuning
To enhance the diversity of the distribution of data generated by the fine-tuned model and maintain better consistency with the distribution of real images [schulman2015trust, schulman2017proximal], denoising diffusion policy optimization (DDPO) [black2023training] is applied in the post-training of OF-Diff. DDPO regards the denoising process of the diffusion model as a multi-step Markov decision process (MDP) (for a detailed derivation, please refer to the Appendix 6.1). To optimize the policy $\pi(a_{t}\mid s_{t})$ so as to maximize the cumulative reward $\mathbb{E}_{\tau\sim p(,\cdot\mid\pi)}!\Bigl[\sum_{t=0}^{T}R(\mathbf{s}_{t},\mathbf{a}_{t})\Bigr]$, the gradient $\hat{g}$ is computed as follows:
$\hat{g}=\mathbb{E}\!\biggl[\sum_{t=0}^{T}\frac{p_{\theta}(\mathbf{x}_{t-1}\mid c,\ t,\ \mathbf{x}_{t})}{p_{\theta^{\prime}}(\mathbf{x}_{t-1}\mid c,\ t,\ \mathbf{x}_{t})}\cdot r(\mathbf{x}_{0},c)\cdot\nabla_{\theta}\log p_{\theta}(\mathbf{x}_{t-1}\mid c,\ t,\ \mathbf{x}_{t})\biggr]$ (8)
$r(\mathbf{x}_{0},c)=\bigl(KNN(\mathbf{x}_{0},\mathbf{x}_{0})-\omega KL(\mathbf{x}_{0},\mathbf{x}_{0}^{\prime})\bigr)$ (9)
The reward functions based on K-Nearest Neighbor (KNN) and KL divergence are introduced, respectively, to optimize the diversity of generated data and the distribution consistency between generated data and real data. $\omega$ is the weight parameter, and $\mathbf{x}_{0}^{\prime}$ is the real image in the dataset. Following standard practice, we compute the KNN in the low-dimensional embedding space of CLIP’s image encoder. The implementation details are shown in Appendix 6.1.
4 Experiments
4.1 Experimental Settings
Datasets. DIOR-R [cheng2022anchor], the rotated variant of DIOR [li2020object], contains 20 categories annotated with oriented bounding boxes; we follow the official 1:1:2 split for training/validation/testing. DOTA-v1.0 [Xia_2018_CVPR] includes 15 categories featuring dense scenes and small objects. We crop the images from DOTA to $512\times 512$ following MMRotate [zhou2022mmrotate], discarding those without valid objects. HRSC2016 [hrsc2016] is a high‐resolution ship detection dataset with a multi‐level hierarchical taxonomy. We use the finest‐grained level, consisting of 26 detailed ship categories. The experiments of this dataset are reported in Appendix 6.3 Unless otherwise specified, we train the diffusion model on the trainset. For downstream detection, we use the trainset annotations as layout and mix generated samples with the real trainset, and report evaluation results on the testset.
Implementation Details. We train OF-Diff separately on each dataset (DIOR/DOTA), based on the Stable Diffusion 1.5 [Rombach_2022_CVPR] pretrained model. Only the ControlNet and shape feature SD decoder are fine-tuned, while all other modules remain frozen. The weighting coefficient $\lambda$ of the consistency loss is set to 1, the k value in KNN is set to 50, and the weight $\omega$ of the KL divergence is set to 2. Training is performed using the AdamW optimizer with a learning rate of 1e-5. The global batch size is set to 64, and training runs for 100 epochs.
Benchmark Methods. We compare our method with state-of-the-art L2I generation models for both remote sensing (AeroGen [tang2025aerogen], CC-Diff [zhang2024cc]), and natural images (LayoutDiffusion [zheng2023layoutdiffusion], GLIGEN [li2023gligen]). For a fair comparison, all models are re-trained using our dataset settings, following their official training details respectively.
Evaluation Metrics. To more comprehensively evaluate the effectiveness of OF-Diff, we adopt a total of 13 metrics spanning 4 different evaluation aspects.
  • Generation Fidelity. We use FID [heusel2017gans] and KID [binkowski2018demystifying] to assess perceptual quality, along with CMMD [jayasumana2024rethinking], which measures CLIP feature distances between generated and real images to evaluate layout alignment.
  • Layout Consistency. We report CAS [ravuri2019classification] using a pretrained classifier to assess object recognizability, and YOLOScore by applying a pretrained Oriented R-CNN [Xie_2021_ICCV] (w/. Swin Transformer backbone [liu2021swin], MMRotate) to generated images for instance-level consistency.
  • Shape Fidelity. To assess the geometric quality of generated instances, we perform pairwise comparisons with ground-truth shapes. Each instance pair is cropped, resized to $\times64$, and converted to edge maps. We compute five metrics: IoU, Dice, Chamfer Distance (CD), Hausdorff Distance (HD), and SSIM [wang2004image].
[…]
Figure 4: Qualitative results on DIOR, DOTA and HRSC2016. OF-Diff is more realistic and fidelity compared to other methods.
4.2 Qualitative Results
Comparative Results. Figure 4 compares the generation results of OF-Diff with other methods. OF-Diff not only generates more realistic images but also has the best controllability. For instance, in the first two cases, OF-Diff successfully controlled the number and layout information of the generated objects. The third and fourth cases demonstrate the accuracy of OF-Diff in generating small targets, which other algorithms fail to do accurately. The last case shows the superiority of OF-Diff over other algorithms when generating objects with complex shapes such as airplane.
Diversity Results. The images generated by OF-Diff consistently present plausible textures and realistic object shapes, as shown in Figure 6 in Appendix. For instance, airplanes rendered at different orientations maintain coherent semantic relationships with their surrounding environments. Even in small-object scenes (some of which are grayscale remote-sensing images from the DOTA dataset), OF-Diff can still generate visually faithful and geometrically accurate results.
Table 1: Quantitative comparison with SOTA methods on DIOR and DOTA. We evaluate performance on generation fidelity (FID, KID, CMMD), layout consistency (CAS, YOLOScore) and trainability (mAP). OF-Diff demonstrates superior overall performance.
  Method DIOR Dataset DOTA Dataset
[…]
LayoutDiff 37.60 0.015 0.447 70.32 7.01 52.14 21.73 0.015 0.288 77.56 21.43 66.75
GLIGEN 35.06 0.010 0.622 76.41 6.51 51.27 39.79 0.026 0.357 76.19 15.58 66.10
AeroGen 27.78 0.013 0.563 81.69 55.38 53.37 26.65 0.017 0.298 81.91 44.85 67.09
$C$-Diff 49.62 0.024 0.685 82.61 42.17 53.48 32.40 0.019 0.279 81.63 49.62 66.52
Ours 24.92 0.011 0.312 82.55 58.99 54.44 20.84 0.014 0.271 83.79 55.68 67.89
Figure 5: AP50 on DIOR and DOTA.
Table 2: Object-Shape Fidelity on Canny Edge Maps. We measure the morphological similarity between generated and ground-truth instances by computing IoU, DICE, Chamfer Distance (CD), Hausdorff Distance (HD), and SSIM.
  Method DIOR Dataset DOTA Dataset
[…]
LayoutDiff 0.0497 0.0908 12.037 25.962 0.1667 0.0402 0.0748 15.229 30.202 0.2194
GLIGEN 0.055 0.1002 12.257 25.850 0.1652 0.0645 0.1182 10.432 23.196 0.1967
AeroGen 0.0855 0.153 8.209 20.314 0.2142 0.0863 0.1536 8.1386 20.687 0.2261
$C$-Diff 0.0891 0.1582 8.0909 20.066 0.1963 0.0692 0.1255 9.6226 21.247 0.2171
Ours 0.1009 0.1763 7.6579 19.459 0.2691 0.1205 0.2045 6.6317 17.311 0.2938
Table 3: Quantitative comparison on the unknown layout dataset during training (DIOR Val).
  Method Unknown Layout during Training
[…]
LayoutDiff 44.58 0.018 0.539 29.34 10.37 30.41 53.07 32.07
GLIGEN 39.56 0.013 0.444 66.36 2.13 30.06 52.68 31.29
AeroGen 28.62 0.013 0.276 80.78 46.36 32.98 55.11 34.26
$C$-Diff 49.92 0.024 0.513 78.01 51.74 32.49 53.72 35.39
Ours 24.18 0.012 0.271 83.34 49.59 33.02 56.65 36.17
Table 4: Ablation study: impact of ESGM, Online-distillation $L_{c}$, and DDPO on semantic consistency (CAS) and downstream trainability (YOLOScore and mAP50).
  ESGM $L_{c}$ DDPO FID $\downarrow$ KID $\downarrow$ CMMD $\downarrow$ CAS $\uparrow$ YOLOScore $\uparrow$ mAP50 $\uparrow$
42.59 0.029 0.965 80.27 41.20 52.13
24.87 0.012 0.428 82.16 55.08 52.76
36.25 0.021 0.596 81.57 46.27 53.14
41.26 0.027 0.815 81.06 42.53 53.41
24.98 0.010 0.313 82.30 57.83 54.31
25.78 0.013 0.368 82.37 58.26 54.17
37.98 0.025 0.613 81.91 47.74 53.21
24.92 0.011 0.312 82.55 58.99 54.44
4.3 Quantitative Results
Generation Fidelity and Consistency. We compared OF-Diff with state-of-the-art generation methods in remote sensing, including layoutDiffusion [zheng2023layoutdiffusion], GLIGEN [li2023gligen], AeroGen [tang2025aerogen], and CC-Diff [zhang2024cc]. The performance of these methods is reported in Table 1. OF-Diff achieved nearly the best performance in both generation fidelity metrics (FID, KID, CMMD) and layout consistency metrics, especially on the DOTA dataset. Additional results are available in the appendix 6.3 for the HRSC2016 dataset.
Trainability of Object Detection. Following the data enhancement protocol in [Chen_Xie_Chen_Wang_Hong_Li_Yeung_2023], we double the training samples using OF-Diff and assess detection results with the expanded dataset. As shown in Table 9 in Appendix 6.6, OF-Diff performs the best on both DIOR and DOTA with mAP improved by 2.2% and 1.94% compared to baseline,respectively. Notably, the performance of several polymorphic and small object classes shows significant improvement. According to Figure 5 (a) and (b), the AP50 increases by 8.3%, 7.7%, and 4.0% for airplane, ship, and vehicle on DIOR, and 7.1%, 5.9% and 4.4% for swimming pool, small vehicle, and large vehicle on DOTA.
Object-Shape Fidelity. We measure the morphological similarity between the generated instances and ground truth by calculating the Intersection over Union (IoU), DICE coefficient, Chamfer distance (CD), Hausdorff distance (HD), and Structural Similarity Index (SSIM), based on the Canny Edge Map. As shown in Table 2, the results demonstrate that OF-Diff attains state-of-the-art performance in all evaluation metrics for object-shape fidelity. Specifically, we first convert the rotated bounding box (R-Box) to a horizontal bounding box (H-Box) and crop the instance with a 20% padding to ensure the full object is captured. The cropped patches are then resized to $64\times 64$ pixels, and their shapes are extracted using cv2.Canny. For a detailed qualitative comparison, Figure 11 in Appendix 6.7 visualizes the instance patches and their corresponding edge maps from different methods. Each image set is ordered as follows: Ground Truth, OF-Diff, AeroGen, CC-Diff, GLIGEN, and LayoutDiff, demonstrating our method’s superior ability to adhere to object shapes.
Adaptability of Unknown Layout. To evaluate robustness of these methods, we also generate images based on the unknown layouts during the training phase. According to Table 3, for unknown layout, OF-Diff performs well in terms of generation fidelity, layout consistency, and trainability. In downstream tasks, OF-Diff still delivers a 1.54% mAP gain over the second-best method.
The Detailed Results of Downstream. Table 10 and 11 in Appendix 6.7 report the average precision (AP) obtained by the competing generative methods over multiple categories in the downstream tasks. From Tables 10, it can be observed that OF-Diff (ours) achieves a clear advantage in several categories. For instance, OF-Diff achieves superior performance on Airplane (71.3%), Golf Field (75.4%) and Ship (70.5%), with improvements of approximately 5% to 10% over the second-best method. For a few other categories, OF-Diff does not deliver the top AP, yet the gap to the best result remains marginal. Table 11 shows that, on the DOTA dataset, OF-Diff obtains the highest AP in roughly half of all categories and still delivers notable gains in categories such as Small-vehicle (68.3%), Ship (84.4%) and Swimming-pool (67.9%).
4.4 Ablation Study
We assessed the impact of different modules on image generation semantic consistency and downstream trainability by OF-Diff in Table 4. We found that the images generated with captions are more in line with semantic consistency and human aesthetics, but the fidelity of these images decreases. This is equivalent to the data distribution deviating from the real dataset and being more inclined towards the data distribution during pre-training. We conduct human/GPT assessments and a fine-grained feature analysis in Appendix 6.6, which collectively reveal the nature of this trade-off. Therefore, the ablation experiments for each module were conducted based on the absence of caption input. The contribution of each module to the enhancement of image generation fidelity is evaluated by incorporating additional components into the diffusion model with online-distillation. DDPO indicates whether to fine-tune the trained diffusion model through reinforcement learning. Results show that Enhanced Shape Generation Module (ESGM), Online-Distillation ($L_{c}$) and the DDPO based on KNN and KL Divergence effectively improve the performance metrics. Notably, ESGM can substantially improve the YOLOScore by over 10%. In addition, we vary the weighting coefficient $\lambda$ in the consistency loss (Eq. 7) to assess its impact on mAP and FID. As shown in Figure 5 (c) and (d), both metrics are optimal at $\lambda=1$.
4.5 Discussion
As shown in Figure 10 in Appendix 6.2, the inclusion of additional captions as input has a significant impact on the outcomes of image generation. Specifically, incorporating captions enhances the aesthetic appeal of the generated images, resulting in richer and more visually pleasing color compositions. However, this improvement comes at a cost: similar to CC-Diff, it leads to a deviation of the generated data distribution from that of the original real data. In contrast, when no additional captions are provided as input, although the generated images may appear less aesthetically refined, their data distribution remains closer to that of real images. A user study from both human and GPT-5 in Table 8 have confirmed this. Further analysis of the generated-image distribution and the impact of aesthetics on performance is provided in the Appendix 6.6.
4.6 Limitations
The proposed OF-Diff injects object shape masks extracted from the image layout as controllable conditions into the diffusion model, which effectively enhances object fidelity and improves the generation of small objects. However, this also makes the model heavily dependent on the quality of the extracted shape masks. If the obtained masks are already distorted, the resulting generated images will likewise fail to achieve satisfactory quality. Therefore, in some scenarios, it is necessary to adopt more suitable mask extraction models (e.g., RemoteSAM, SAM2, SAM3) to obtain higher-quality masks. In future work, we will explore stronger object mask extraction and generation strategies to better ensure object-level generation quality while improving the texture details of the synthesized images. We will also investigate more advanced models and training paradigms to further promote the development of remote-sensing image generation.
5 Conclusion
Existing image generation methods struggle to precisely generate dense small objects and those with complex shapes, such as numerous small vehicles and airplanes in remote sensing images. To address this, we introduce OF-Diff, an online-distillation controllable diffusion model with prior shapes extraction. During the training phase, we extract the prior shape of objects to enhance controllability and use a online-distillation diffusion with parameter sharing to improve the model’s learning ability for real images. Therefore, in the sampling phase, OF-Diff can generate images with high fidelity without real images as references. Finally, we fine-tune the diffusion by DDPO that combines KNN and KL divergence to make the synthesized images more realistic and consistent. Extensive experiments demonstrate the effectiveness and superiority of OF-Diff in generating small and difficult objects with complex structures and dense scenes in remote sensing.